OpenStudy (anonymous):

On a certain day, the Wilton County Jail had 183 prisoners. Of these, 127 were accused of felonies, and 117 were accused of misdemeanors. How many prisoners were accused of both a felony and a misdemeanor

OpenStudy (anonymous):

 Let k be the number of prisoners convicted for both felony and misdemeanor. x = 127 - k y = 117 - k Total prisoners = x + y + k = 183 (127 - k ) + k + (117 - k) = 183 244 - k = 183 k = 61